A sample of gas at 25.0oC has a volume of 3.20L at a pressure of 100kPa. What would its volume be if the pressure was increased to 1.3atm and the temperature was decreased to -57.0oC?
1
Expert's answer
2020-10-12T13:29:46-0400

T1=25.0 C= 298 K

V1=3.2 L

p1=100 kPa

T2=-57.0 C = 216 K

p2=1.3 atm= 101.3 * 1.3 (kPa)=131.7 kPa

V2 - ?



(p1V1)/T1=(p2V2)/T2;

V2=(p1V1*T2)/(T1*p2);

V2=(100*3.2*216)/(298*131.7)=1.76 L

Answer : 1.76 (L)